OpenStudy (anonymous):

Helppp please! Find the solutions to the equation. 6x^2+7x-5=0

OpenStudy (anonymous):

Use the quadratic equation on one like this.

OpenStudy (amistre64):

10-3 = 7.... which leads me to believe this can be done easilier :)

OpenStudy (anonymous):

maybe it factors nicely?

OpenStudy (anonymous):

ah, (2x-1)(3x+5)

OpenStudy (amistre64):

6(-5) = -30 10(-3) = -30; 10-3=+7 (x+10/6) (x-3/6) ; simplify

OpenStudy (amistre64):

(x+5/3) (x-1/2) ; if its a fraction stick the bottom in front (3x+5)(2x-1)

OpenStudy (anonymous):

okay so then i foil correct?

OpenStudy (amistre64):

"foil" is a four letter word :) foil is simply a memory device used to teach people how to multiply ...

OpenStudy (anonymous):

\[(-b \pm \sqrt{b ^{2}-4ac})\div2a\]

OpenStudy (amistre64):

the hardest thing about factoring is that you have to predict the future to some extent

jhonyy9 (jhonyy9):

-7 +/- 13 x_1,_2 = ---------- = -20/12 and 6/12 12

OpenStudy (amistre64):

to find out what x =; we can stop at the fraction part and study the factors

jhonyy9 (jhonyy9):

what you can simplifie again

OpenStudy (amistre64):

(x+5/3) (x-1/2) = 0 means that either x+5/3 = 0 or x-1/2 = 0 solve each one for 0

OpenStudy (anonymous):

I got it ! it's x=1/2 and x=-5/3

OpenStudy (amistre64):

x = -5/3 or x=1/2 very good

OpenStudy (anonymous):

Thank you!